There are 4 times as many boys as girls in a certain English class. If there are n girls, how many boys are there? how many students in all? State by an equation that there are 30 students in a class. Find the number of boys
Answer by ewatrrr(24785)


Hi,
4 times as many boys as girls
Let n and 4n represent the number of girls and boys respectively
Question states***
n + 4n = 30
5n = 30
n = 6, the number of girls. Number of boys = 24 (4*6)
CHECKING our Answer***
6 + 24 = 30